If the output variable y is 38 less than the input variable x, what is the input corresponding to an output value of 121?
I thought I was starting to understand this, I was wrong.
Thanks for your help!!

Found 2 solutions by Leaf W., solver91311:
Answer by Leaf W.(135)

"the output variable y is 38 less than the input variable x"
Therefore, y is 38 less than x. Since "is" is a signal word for "equals," and 38 less than x is the same thing as x minus 38, you can express this as y = x - 38
Next, since y is defined as the output, and the problem tells that the output value is 121, plug in 121 for y.
121 = x - 38
Since the problem asks for the input, which it also states to be x ("the input variable x"), solve for x to find the input by adding 38 to both sides.
121 + 38 = x - 38 + 38
Simplify by adding 121 and 38. -38 and 38 cancel each other out.
159 = x
==> Therefore, the input value, x, is 159 when the output value, y, is 121.

Answer by solver91311(24713)



Since there is a constant difference between the input and output it is a linear relationship. So y is always 38 less than x. So , and so .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism